The frequency of the radio wave emitted by a cordless telephone is 980 MHz. What is the wavelength of that wave? λ = Answer in m
Answer #1

Frequency * wavelength = speed of wave

As we know that speed of radio wave is 3*10^8 m/s

So wavlength. = 980*10^6/3*10^8 = 980/300 = 3.26666 m